OpenStudy (anonymous):

I NEED HELP PLEASE f(x)=5x-x^2, f(5+h)-f(5) all over (h) h cannot = 0 go over step by step so i know how to do this

OpenStudy (anonymous):



Nnesha (nnesha):

first find f(5+h) f(5+h) means x = 5+h so substitute x for 5+h into f(x) function

Nnesha (nnesha):

\[\huge\rm f(\color{ReD}{x})=5\color{ReD}{x}-\color{ReD}{x}^2\] to find f(5+h) x with 5+h \[\huge\rm f(\color{ReD}{5+h})=5(\color{ReD}{5+h})-(\color{ReD}{5+h})^2\] distribute and then combine like terms to find f(5) replace x with 5